# Orlicz-Sobolev Transport for Unbalanced Measures on a Graph

### v1

@@ 3.1 Orlicz-EPT: EPT on a Graph with Orlicz Structure
We next show that the objective function of Orlicz-EPT w.r.t. $t$ is monotonically non-increasing.
@@
For unbalanced measures $\mu,\nu\in{\mathcal{P}}({\mathbb{G}})$, construct corresponding $\hat{\mu},\hat{\nu},\hat{c}$. For an $N$-function $\Phi$, let
Then, $\mathcal{A}(t;\hat{\mu},\hat{\nu})$ is monotonically non-increasing w.r.t. $t$.
Proof is placed in Appendix §A.1.
@@
Observe that for a fixed $t$, $\mathcal{A}$ is a standard OT problem between $\hat{\mu}$ and $\hat{\nu}$ with the cost function $\Phi\!\left(\frac{\hat{c}(\cdot,\cdot)}{t}\right)$. We show that the monotonicity is preserved for its corresponding entropic regularization.
@@
$\mathcal{A}_{\varepsilon}(t;\hat{\mu},\hat{\nu})\!:=\!\inf_{\tilde{\gamma}\in\Pi(\hat{\mu},\hat{\nu})}\!\left[\int_{\hat{\mathbb{G}}\times\hat{\mathbb{G}}}\Phi\left(\!\frac{\hat{c}(x,y)}{t}\!\right)\!\mathrm{d}\tilde{\gamma}(x,y)\!-\!\varepsilon H(\tilde{\gamma})\right],$ (11)
where $\varepsilon\geq 0$ and $H$ is Shannon entropy defined as $H(\tilde{\gamma}):=-\int_{\hat{\mathbb{G}}\times\hat{\mathbb{G}}}(\log\tilde{\gamma}(x,y)-1)\mathrm{d}\tilde{\gamma}(x,y)$. Then, $\mathcal{A}_{\varepsilon}(t;\hat{\mu},\hat{\nu})$ is monotonically non-increasing w.r.t. $t$.
Proof is placed in Appendix §A.2.
Additionally, we derive the upper and lower bounds for $\mathcal{A}_{\varepsilon}$.
@@
We have
$\mathcal{A}_{\varepsilon}\!\left(\frac{\mathcal{W}_{\hat{c}}(\hat{\mu},\hat{\nu})}{\Phi^{-1}(1+\varepsilon\left[H(\hat{\mu})+H(\hat{\nu})-1\right])};\hat{\mu},\hat{\nu}\right)\geq 1,$
$\mathcal{A}_{\varepsilon}\!\left(\frac{L_{\hat{\mu},\hat{\nu}}}{\Phi^{-1}(1+\varepsilon)};\hat{\mu},\hat{\nu}\right)\leq 1,$
where $L_{\hat{\mu},\hat{\nu}}=\max_{x\in\mathrm{supp}(\hat{\mu}),y\in\mathrm{supp}(\hat{\nu})}\hat{c}(x,y)$, and $\mathrm{supp}(\cdot)$ is a set of supports of a measure.
Proof is placed in Appendix §A.3.
From the monotonicity of $\mathcal{A}_{\varepsilon}$ in Proposition 3.3, and the limits of $\mathcal{A}_{\varepsilon}$ in Proposition 3.4, we can leverage the binary search to compute the entropic regularized Orlicz-EPT (corresponding to the original Orlicz-EPT (9)), defined as:
@@
Although the entropic regularized Orlicz-EPT $\mathcal{OE}_{\Phi,\varepsilon}$ (12) scales up the computation of the original Orlicz-EPT $\mathcal{OE}_{\Phi}$ (9) by leveraging binary search on the quadratic-complexity $\mathcal{A}_{\varepsilon}$ (11) instead of the super-cubic-complexity $\mathcal{A}$ (10), the nature two-level optimization complexity of $\mathcal{OE}_{\Phi,\varepsilon}$ still remains, which hinders its practical applications, especially in large-scale settings. To tackle this computational challenge, we next leverage a collection of special convex functions (i.e., $N$-functions) to propose the Orlicz-Sobolev transport, which can adopt Orlicz geometric structure in the same sense as the Orlicz-EPT, and especially, it is much more efficient for computation.
3.2 Orlicz-Sobolev Transport
We recall the dual formula of EPT, and the graph-based Orlicz-Sobolev space which play as the key components for our proposed Orlicz-Sobolev transport (OST) for unbalanced measures on a graph.
Dual formula of EPT on a graph (Le et al., 2023).
Assume that $\lambda\geq 0$ and the nonnegative weight functions $w_{1},w_{2}$ are $b$-Lipschitz w.r.t. $d_{\mathbb{G}}$. Following Le et al. (2023, Corollary 3.2), the dual EPT on a graph is as follows:
@@ 3.2 Orlicz-Sobolev Transport
Graph-based Orlicz-Sobolev space (Le et al., 2024).
Such function $h$ is unique in $L_{\Phi}({\mathbb{G}},\omega)$ and is called the generalized graph derivative of $f$ w.r.t. the measure $\omega$. Henceforth, this generalized graph derivative of $f$ is denoted $f^{\prime}$.
Orlicz-Sobolev transport (OST).
Inspired by the generalized Sobolev transport (Le et al., 2024), we exploit the dual EPT on a graph. We then replace the Lipschitz constraint for the critic function $f$ in $\mathbb{U}$ by a constraint involving the graph-based Orlicz-Sobolev space, i.e., $f\in W\!L_{\Psi}({\mathbb{G}},\omega),\left\lVert f^{\prime}\right\rVert_{L_{\Psi}}\leq b$ where $\Psi$ is the complement $N$-function of $\Phi$. Moreover, thanks to the observation in Le & Nguyen (2021), from Equation (14) and $\left\lVert f^{\prime}\right\rVert_{L_{\Psi}}\leq b$, we employ the generalized Hölder inequality w.r.t. Luxemburg norm in Orlicz space $L_{\Psi}({\mathbb{G}},\omega)$ (Adams & Fournier, 2003, §8.11), then $\int_{[z_{0},x]}f^{\prime}(y)\omega(\mathrm{d}y)$ is controlled by $2b/\Phi^{-1}\!\left(\omega([z_{0},x])^{-1}\right)$. Therefore, instead of the bounded constraint on the critic function $f$ in $\mathbb{U}$, we constraint only on $f(z_{0})$.
Definition 3.5 (Orlicz-Sobolev transport).
Let $\omega$ be a nonnegative Borel measure on graph ${\mathbb{G}}$, and ${\mathcal{I}}_{\alpha}:=\left[-w_{2}(z_{0})-\frac{b\lambda}{2}+\alpha,w_{1}(z_{0})+\frac{b\lambda}{2}-\alpha\right]$ where $\alpha\in[0,\frac{1}{2}(b\lambda+w_{1}(z_{0})+w_{2}(z_{0}))]$. Given a pair of complementary $N$-functions $\Phi,\Psi$, [^2] We give a review of complementary $N$-functions in §B.3. and for $\mu,\nu\in{\mathcal{P}}({\mathbb{G}})$, the Orlicz-Sobolev transport is defined as follows
@@
where $\mathbb{U}_{\Psi,\alpha}\!:=\!\big{\{}f\!\in\!W\!L_{\Psi}({\mathbb{G}},\omega),\left\lVert f^{\prime}\right\rVert_{L_{\Psi}}\!\leq\!b,f(z_{0})\!\in\!{\mathcal{I}}_{\alpha}\big{\}}$.
Intuitively, $\mathbb{U}_{\Psi,\alpha}$ is the collection of all functions $f$, expressed by $f(x)=s+\int_{[z_{0},x]}h(y)\omega(\mathrm{d}y),\forall x\in{\mathbb{G}}$, where $s\in{\mathcal{I}}_{\alpha}$, and $\left\lVert h\right\rVert_{L_{\Psi}}\leq b$. The upper bound constraint on $\alpha$ is to ensure that ${\mathcal{I}}_{\alpha}$ is nonempty. When $\alpha=0$, the interval ${\mathcal{I}}_{\alpha}$ is the largest. Additionally, OST for unbalanced measures on a graph is an instance of the integral probability metric (Müller, 1997).
@@
We show that OST can be efficiently computed by simply solving a corresponding univariate optimization problem.
Theorem 3.6 (OST as a univariate optimization problem).
Given two unbalanced measures $\mu,\nu\in{\mathcal{P}}({\mathbb{G}})$, define
@@
Then, Orlicz-Sobolev transport can be computed as follows:
@@
The proof is placed in Appendix §A.4.
We next derive the discrete case for the OST. Especially, we provide an explicit expression for the integral in Equation (14).
@@
Let $\omega$ be the length measure of graph ${\mathbb{G}}$, and input measures $\mu,\nu$ are supported on nodes in $V$ of graph ${\mathbb{G}}$ [^3] It can be extended for measures supported in ${\mathbb{G}}$ (see §B.8). Then, we have
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)=\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|+\inf_{k>0}\frac{1}{k}\left(1+\sum_{e\in E}w_{e}\Phi\!\left(kb\left|\mu(\gamma_{e})-\nu(\gamma_{e})\right|\right)\right),$ (20)
The proof is placed in Appendix §A.5.
Therefore, the OST (Definition 3.5) can be efficiently computed by simply solving the univariate optimization problem (20) (e.g., by second-order methods).
@@
Although ${\mathbb{G}}$ is assumed to be a physical graph (§2), Corollary 3.7 implies that OST only depends on the graph structure $(V,E)$ and edge weights $w_{e}$ when measures are supported on nodes in $V$ of graph ${\mathbb{G}}$, i.e., OST is applicable for non-physical graph ${\mathbb{G}}$.
@@
Corollary 3.7 also implies that one can compute OST with $N$-function $\Phi$ without involving its complementary $N$-function $\Psi$ (Equation (20)), unlike its definition (Equation (15)). It requires that $\Psi$ is finited-value to derive the univariate optimization formula (20)), which is satisfied for any $N$-function $\Phi$, i.e., growing faster than linear.
Preprocessing for $\gamma_{e}$. Following the observation in Le et al. (2024), we precompute set $\gamma_{e}$ (Equation (2)) for all edge $e$ in ${\mathbb{G}}$. More concretely, we apply the Dijkstra algorithm to recompute the shortest paths from the root node $z_{0}$ to all other vertices in $V$ with complexity $\mathcal{O}(|E|+|V|\log{|V|})$, where ||$\cdot$|| ⋅ | is the set cardinality.
Remark 3.10 (Sparsity).
Observe that $\forall x\in\text{supp}(\mu)$, its mass is gathered into $\mu(\gamma_{e})$ if and only if $e\subset[z_{0},x]$ (Le et al., 2024). Therefore, by defining set $E_{\mu,\nu}\subset E$ as
$E_{\mu,\nu}:=\left\{e\!\in\!E\mid\exists z\!\in\!(\text{supp}(\mu)\cup\text{supp}(\nu)),e\subset[z_{0},z]\right\}\!,$
it suffices to compute the summation only over edges $e\in E_{\mu,\nu}$ in Equation (20) for OST, i.e., screening all edges $e\in E\setminus E_{\mu,\nu}$.
@@ 4 Theoretical Properties
Orlicz-EPT has its own interesting characteristics. However, it has a high computational complexity due to its two-level optimization formula, which hinders its practical applications. Thus, in this section, we mainly focus on theoretical properties for Orlicz-Sobolev transport (OST). [^4] See §B.8 for Orlicz-EPT.
4.1 Geometric Structure of Orlicz-Sobolev Transport
We derive the geometric structure of OST. Under certain conditions, we prove that OST is a metric.
Proposition 4.1 (Geometric structures of OST).
Let $\omega$ be a nonnegative Borel measure on graph ${\mathbb{G}}$. Assume that $\lambda,w_{1}(z_{0}),w_{2}(z_{0})\geq 0$, and $0\leq\alpha<\frac{b\lambda}{2}+\min\{w_{1}(z_{0}),w_{2}(z_{0})\}$. Then, $\forall\mu,\nu,\sigma\in{\mathcal{P}}({\mathbb{G}})$, we have
@@ 4.1 Geometric Structure of Orlicz-Sobolev Transport
The proof is placed in Appendix §A.6.
4.2 Relations to Other Transport Distances
We draw several connections for OST with other transport distances on graph ${\mathbb{G}}$, including generalized Sobolev transport (GST) (Le et al., 2024), Sobolev transport (ST) (Le et al., 2022), unbalanced Sobolev transport (UST) (Le et al., 2023), Orlicz-EPT (§3.1), a variant of regularized EPT (Le & Nguyen, 2021), and standard optimal transport (OT). Additionally, we consider the limit case $\Phi(t)=t$ for $N$-function, and the special case when graph ${\mathbb{G}}$ is a tree.
Connection with generalized Sobolev transport.
@@ 4.2 Relations to Other Transport Distances
Assume that $\mu({\mathbb{G}})=\nu({\mathbb{G}})$, $b=1$, and denote $\mathcal{GS}_{\Phi}$ for the GST with $N$-function $\Phi$, [^6] See §B.5 for a review on GST $\mathcal{GS}_{\Phi}$. then we have
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)=\mathcal{GS}_{\Phi}(\mu,\nu).$
The proof is placed in Appendix §A.7.
Connection with Sobolev transport.

### v2

@@ 3.2 Orlicz-EPT
It should be noted that, similar to OW, Orlicz-EPT (8) is derived from the standard OT problem (7), which circumvents all challenges coming from the setting of unbalanced measures.
We next show that the objective function of Orlicz-EPT is monotone non-increasing w.r.t. $t$.
@@
Let $\Phi$ be an $N$-function, and let $\hat{c}$ be the cost given by (5). For any probability measures $\hat{\mu}$ and $\hat{\nu}$ on $\hat{\mathbb{G}}$, define
$\mathcal{A}(t;\hat{\mu},\hat{\nu}):=\inf_{\tilde{\gamma}\in\Pi(\hat{\mu},\hat{\nu})}\int_{\hat{\mathbb{G}}\times\hat{\mathbb{G}}}\Phi\!\left(\frac{\hat{c}(x,y)}{t}\right)\mathrm{d}\tilde{\gamma}(x,y)\quad\mbox{for}\quad t>0.$ (9)
[…]
The proof is placed in Appendix §A.2.2.
@@
Observe that for a fixed $t$, $\mathcal{A}$ is a standard OT problem between $\hat{\mu}$ and $\hat{\nu}$ with the cost function $\Phi\!\left(\frac{\hat{c}(\cdot,\cdot)}{t}\right)$. For computational efficiency, [^6] Entropic regularized OT reduces the computational cost of OT from super-cubic to quadratic complexity [20]. we consider its corresponding entropic regularization [20], and show that the monotonicity is preserved.
@@
[…]
The proof is placed in Appendix §A.2.3.
In addition, we obtain the following upper and lower bounds for $\mathcal{A}_{\varepsilon}$.
@@
Let $\mathrm{supp}(\cdot)$ be a set of supports of a measure, then we have
$\mathcal{A}_{\varepsilon}\!\left(\frac{\mathcal{W}_{\hat{c}}(\hat{\mu},\hat{\nu})}{\Phi^{-1}(1+\varepsilon\left[H(\hat{\mu})+H(\hat{\nu})-1\right])};\hat{\mu},\hat{\nu}\right)\geq 1,\quad\text{and}\quad\mathcal{A}_{\varepsilon}\!\left(\frac{L_{\hat{\mu},\hat{\nu}}}{\Phi^{-1}(1+\varepsilon)};\hat{\mu},\hat{\nu}\right)\leq 1,$
where $L_{\hat{\mu},\hat{\nu}}:=\max_{x\in\mathrm{supp}(\hat{\mu}),y\in\mathrm{supp}(\hat{\nu})}\hat{c}(x,y)$.
The proof is placed in Appendix §A.2.4.
Thanks to the monotonicity of $\mathcal{A}_{\varepsilon}$ in Proposition 3.4 and the limits of $\mathcal{A}_{\varepsilon}$ in Proposition 3.5, we can leverage the binary search approach to compute the entropic regularized Orlicz-EPT, which corresponds to the original Orlicz-EPT (8). Precisely, this entropic regularization is defined as
@@
Discussions. Orlicz-EPT is a novel variant of EPT that incorporates Orlicz geometric structure. Leveraging Caffarelli & McCann [12]’s insights and carefully calibrating the ground cost of the corresponding standard OT to ensure its nonnegativity, we are able to bypass all challenges of unbalanced measures and derive the proposed Orlicz-EPT from the standard OT, similar to OW [67]. We note that $\mathcal{OE}_{\Phi,\varepsilon}$ (11) performs binary search with quadratic complexity $\mathcal{A}_{\varepsilon}$ (10), instead of dealing with super-cubic complexity $\mathcal{A}$ (9) in $\mathcal{OE}_{\Phi}$ (8). Unfortunately, the two-level optimization structure of $\mathcal{OE}_{\Phi,\varepsilon}$ still retains significant complexity, severely limiting its practical applications, particularly in large-scale settings. To address this computational challenge, in the next section we exploit the dual EPT and graph structure to develop a novel regularization approach, resulting in the proposed Orlicz-Sobolev transport. This approach adopts the Orlicz geometric structure used in Orlicz-EPT, but offers a much more efficient computation.
4 Orlicz-Sobolev Transport: A Scalable Variant of Orlicz-EPT
In this section, we leverage the dual EPT and underlying graph structure to develop a novel regularization approach, resulting in the proposed Orlicz-Sobolev transport (OST).
Dual EPT. For $b$-Lipschitz $w_{1},w_{2}$ (w.r.t. $d_{\mathbb{G}}$), from [40, Corollary 3.2], the dual EPT is
@@ 4 Orlicz-Sobolev Transport: A Scalable Variant of Orlicz-EPT
Let $\Psi$ be the complement $N$-function of $\Phi$ and $\omega$ be a nonnegative Borel measure on $\mathbb{G}$}. Then let $W\!L_{\Psi}({\mathbb{G}},\omega)$ be the graph-based Orlicz-Sobolev space [41, Definition 3.1] associated to $\Psi$ and $\omega$. Inspired by the approach of GST [41], we consider the critic function $f\in\mathbb{U}$ within $W\!L_{\Psi}({\mathbb{G}},\omega)$. Consequently, the $b$-Lipschitz constraint on the critic function $f\in\mathbb{U}$ is replaced by $\left\lVert f^{\prime}\right\rVert_{L_{\Psi}}\leq b$.
For $f\in W\!L_{\Psi}({\mathbb{G}},\omega)$, we have $f(x)=f(z_{0})+\int_{[z_{0},x]}f^{\prime}(y)\omega(\mathrm{d}y),\forall x\in{\mathbb{G}}$. Let $\mathbf{1}$ be the indicator function. Then by using the generalized Hölder inequality [2, §8.11] and $\left\lVert f^{\prime}\right\rVert_{L_{\Psi}}\leq b$, we can control the integral part over the generalized graph derivative $f^{\prime}$ for $f(x)$ as follows:
$\int_{[z_{0},x]}f^{\prime}(y)\omega(\mathrm{d}y)\leq 2\left\lVert f^{\prime}\right\rVert_{L_{\Psi}}\left\lVert\mathbf{1}_{[z_{0},x]}\right\rVert_{L_{\Phi}}\leq 2b\left\lVert\mathbf{1}_{[z_{0},x]}\right\rVert_{L_{\Phi}}\leq\frac{2b}{\Phi^{-1}\!\left(1/\omega({\mathbb{G}})\right)},$ (13)
where the last inequality is due to the increasing property of $N$-function $\Phi$. Therefore, instead of the bounded constraint on the critic function $f$ in $\mathbb{U}$, we constraint only on $f(z_{0})$.
Definition 4.1 (Orlicz-Sobolev transport (OST)).
For $\alpha\in[0,\frac{1}{2}(b\lambda+w_{1}(z_{0})+w_{2}(z_{0}))]$, let ${\mathcal{I}}_{\alpha}:=\left[-w_{2}(z_{0})-\frac{b\lambda}{2}+\alpha,w_{1}(z_{0})+\frac{b\lambda}{2}-\alpha\right]$. The Orlicz-Sobolev transport for $\mu,\nu\in{\mathcal{P}}({\mathbb{G}})$ is defined
@@
where $\mathbb{U}_{\Psi,\alpha}:=\big\{f\in W\!L_{\Psi}({\mathbb{G}},\omega):\,\left\lVert f^{\prime}\right\rVert_{L_{\Psi}}\leq b,f(z_{0})\in{\mathcal{I}}_{\alpha}\big\}$.
Intuitively, $\mathbb{U}_{\Psi,\alpha}$ is the collection of all functions $f$ expressed by $f(x)=s+\int_{[z_{0},x]}h(y)\omega(\mathrm{d}y),\forall x\in{\mathbb{G}}$, where $s\in{\mathcal{I}}_{\alpha}$, and $\left\lVert h\right\rVert_{L_{\Psi}}\leq b$. The upper bound constraint on $\alpha$ is to ensure that ${\mathcal{I}}_{\alpha}$ is nonempty. When $\alpha=0$, ${\mathcal{I}}_{\alpha}$ is the largest interval. Also, OST is an instance of the integral probability metric [49].
@@
Given unbalanced measures $\mu,\nu\in{\mathcal{P}}({\mathbb{G}})$, for brevity let us define
@@
Theorem 4.2 (Univariate optimization problem for OST).
OST can be computed as follows
@@
The proof is placed in Appendix §A.2.5.
We derive the discrete case for OST which provides an explicit expression for the integral in (65).
@@
Let $\omega$ be the length measure of graph $\mathbb{G}$}, and assume that input measures $\mu,\nu$ are supported on nodes in $V$ of graph $\mathbb{G}$}. [^7] It can be extended for measures supported in $\mathbb{G}$} (see §B.2). Then, we have
$\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)=\Theta|\mu({\mathbb{G}})-\nu({\mathbb{G}})|+\inf_{k>0}\frac{1}{k}\left(1+\sum_{e\in E}w_{e}\Phi\!\left(kb\left|\mu(\gamma_{e})-\nu(\gamma_{e})\right|\right)\right).$ (19)
The proof is placed in Appendix §A.2.6.
Therefore, OST can be efficiently computed by simply solving the univariate optimization problem (19), thanks to the proposed novel regularization for critic functions in $\mathbb{U}_{\Psi,\alpha}$.
@@
In §2, $\mathbb{G}$} is assumed to be a physical graph. Corollary 4.3 implies that OST only depends on graph structure $(V,E)$ and edge weights $w_{e}$ when input measures are supported on nodes in $V$ of $\mathbb{G}$}. Hence, OST is applicable for non-physical graph $\mathbb{G}$} for such cases.
@@
Corollary 4.3 also implies that one can compute OST with $N$-function $\Phi$ without involving its complementary $N$-function $\Psi$ (19), unlike its definition (14). The univariate optimization formula (19)) for OST requires that $\Psi$ is finite-valued, which is satisfied for any NN-function $\Phi$ as it grows faster than linear.
Preprocessing for $\gamma_{e}$. Similar to the GST computation [41], we precompute set $\gamma_{e}$ (1) for all edge $e$ in $\mathbb{G}$}. More concretely, we apply the Dijkstra algorithm to recompute the shortest paths from $z_{0}$ to all other vertices in $V$ with complexity $\mathcal{O}(|E|+|V|\log{|V|})$, where $|\cdot|$ denotes the set cardinality.
Sparsity. Observe that for every $x\in\text{supp}(\mu)$, its mass is gathered into $\mu(\gamma_{e})$ if and only if $e\subset[z_{0},x]$ [41]. Let $E_{\mu,\nu}\hskip-1.99997pt:=\hskip-1.99997pt\left\{e\!\in\!E\mid\exists z\!\in\!(\text{supp}(\mu)\cup\text{supp}(\nu)),e\subset[z_{0},z]\right\}\subset E$. Then it suffices to compute the summation only over edges $e\in E_{\mu,\nu}$ in (19) for OST, i.e., screen out all edges $e\in E\setminus E_{\mu,\nu}$.
@@ 5 Theoretical Properties
In this section, we leverage the computational efficiency of OST to derive its geometric structure and explore its connections to other transport distances.
Geometric structures of OST.
Proposition 5.1 (Geometric structure).
Let $0\leq\alpha<\frac{b\lambda}{2}+\min\{w_{1}(z_{0}),w_{2}(z_{0})\}$ and $\mu,\nu,\sigma\in{\mathcal{P}}({\mathbb{G}})$.
@@
We next establish connections of OST with other transport distances, including GST [41], Sobolev transport (ST) [39], unbalanced Sobolev transport (UST) [40].
Connection of OST with GST. Denote $\mathcal{GS}_{\Phi}$ for the GST with $N$-function $\Phi$.
@@
For $\mu({\mathbb{G}})=\nu({\mathbb{G}})$, $b=1$, then $\mathcal{OS}_{\Phi,\alpha}(\mu,\nu)=\mathcal{GS}_{\Phi}(\mu,\nu)$.
Connection of OST with ST. Denote $\mathcal{S}_{p}$ for the $p$-order ST, for $1<p<\infty$.